A 0.260 kg ball approaches a bat horizontally with a speed of 18.0 m/s and after getting hit by the bat, the ball moves in the opposite direction with a speed of 22 m/s. If the ball is in contact with the bat for 0.0600 s, determine the magnitude of the average force exerted on the bat.
1
Expert's answer
2020-12-04T09:35:55-0500

We can find the average force exerted on the bat from the impluse-momentum change equation:


"F\Delta t=m\Delta v,""F=\dfrac{m\Delta v}{\Delta t},""F=\dfrac{0.260\ kg\cdot(-22\ \dfrac{m}{s}-18\ \dfrac{m}{s})}{0.06\ s}=-173.3\ N."


The magnitude of the average force exerted on the bat is 173.3 N. The sign minus means that the average force exerted on the bat directed away from the bat.

Answer:

"F=173.3\ N," away from the bat.